Seawater Maintenance Log
 Share
The version of the browser you are using is no longer supported. Please upgrade to a supported browser.Dismiss

 
View only
 
 
ABCDEFGHIJKLMNOPQRST
1
Date (yyyymmdd)Time (PST)Task
2
20181017Cleaned fluoro and trans
3
20181012Cleaned fluor, trans, SUNA
4
20180920Cleaned fluoro and trans
5
20180827Cleaned fluoro and transmission. Sent pCO2 sensor back to Turner for inspection. I suspect Durafet electrode has failed.
6
20180824pCO2 readings are not believed to be accurate. Sensor placed in bucket of shallow freshwater to equilibrate with ambient pCO2. It stablized around 1200ppm over the weekend.
7
20180824pH probe flatlined after severa days of highly erratic output.
8
20180821Calibrated pH with a Dickson sample.
9
20180820Installed new durafet ph electrode
10
201808021200Cleaned, SUNA, fluoro, trans, optode, CTD, pCO2
11
20180619930Cleaned fluoro and trans
12
201806111130Cleaned fluor and trans..
13
201805311100Cleaned fluor and trans. Data seem spikey. Keeping an eye on it.
14
20180523CTD not logging properly. collected pH sample at noon local time. Calibrated pH sensor with Dickson standard. Took another pH sample afterward.
15
201805221200Cleaned fluoro and trans, pCO2, optode, pH. Swapped CTD. Pump issues occured. Facilities team are trying to reprime pumps.
16
201805111430Cleaned fluoro and trans
17
201805041130Cleaned fluoro and trans
18
201804271400pCO2 sensor back from manufacturer service.. deployed
19
201804261400Cleaned fluoro and trans
20
20180404930Cleaned fluoro, trans, CTD. Removed pCO2 sensor to return to manufacturer for annual maintenance.
21
201804031300Cleaned fluoro and trans
22
20180327815Cleaned fluoro and trans
23
20180322715Swapped optode 1238 for 1921. Optode 1921 seemed to be reporting higher than normal values. Cleaned trans
24
201803211100cleaned fluoro and trans
25
20180315900cleaned fluoro and trans
26
201803091200cleaned fluoro and trans
27
2018030514:00Cleaned fluoro, trans, nitrate, optode, pco2, CTD sensors.
28
201803021100Cleaned fluoro and trans
29
201802211000Cleaned fluoro and trans
30
20180210-13CTD memory full. No data collected.
31
201802081300Intake Screens cleaned, zincs installed
32
20180206900Cleaned fluoro, trans, CTD, and pCO2 sensors.
33
2018020112:00Intake pipes cleaned resulting in noticable change in FLUORO data.
34
201801241400Cleaned trans, fluoro, and SUNA
35
201801231606Collected grab sample for Nitrate analysis.
36
201801231142Collected grab sample for Nitrate analysis. Cleaned fluoro and trans.
37
2018011822:10Pump 2 lost prime resulting in inadequate flow to sensors until 8:30 the next morning.
38
201801161510Swapped DO optode (1921 for 1238).. will send back for service. Cleaned fluoro and trans.
39
201801121400cleaned fluoro and trans
40
201801091145Cleaned fluoro and trans
41
20180104Deployed Satlantic SUNA Nitrate sensor in pumphouse. Cleaned fluoro and trans.
42
2018-01-021100Cleaned fluoro and trans
43
20171224 to 27Suspect inadequate flow to sensors
44
2017-12-21-22Power outage at Lab resulting in some data loss
45
2017-12-2110:00Cleaned fluoro and trans
46
20171219-20Data loss CTD not responding.. IT issues.
47
20171215-16Bad data.. likely due to clog in the flowline.. cleared itself over the weekend
48
20171215930Cleaned fluoro, trans, and pCO2 sensors
49
201712121030Cleaned fluoro and trans
50
20171203 to 04Suspect poor flow to sensors
51
20171201 to 02Suspect poor flow to sensors
52
201711301300Cleaned fluoro, trans, pH, CTD, and optode. Existing KCL gel came out of electrode tip. replace with some expired gel, but data are not being reported correctly. Will have to order new gel?
53
201711291100Cleaned SWI fluorometer and transmissometer. Ordered new tubing for discharge.
54
20171013-20171116Jason Adelaars gone on vacation. NOAA tide server discontinued, causing fouling of whole data acquisition program resulting in loss of data. New tide API implemented on 20171129. Flaw in data acquisition corrected.
55
201710051200Fluorometer and Transmissometer returned with fresh calibration. Deployed today. Trained Tom & Cameron on how to clean optical sensors.
56
20170914morningDived and cleaned the intake pipes & screens. Lots of growth on the screens, so brushed them thoroughly.
57
20170911All day Noticed flow issues in the data this morning. Cleared biofouling and cleaned: pCO2, CTD, DO, and pH probes
58
20170829Some data lost during updating of pCO2 data processing. Updated pCO2 to correct for total dissolved gas pressure in water
59
20170811pCO2 data back to real-time!
60
20170718 evening to next morningTubing to pH sensor came detached. No flow to pH sensor. All other sensors recieveing ok flow.
61
2017071812:00Cleaned pCO2, CTD, and pH sensor. Took transmissometer and fluorometer offline for service and calibration.
62
20170713-17Loss of some data while merging pCO2 dataset with the rest of the seawater database.
63
2017071011:00Cleaned fluoro and trans
64
20170705-07Server down for a couple days. Couldn't log in due to lockout policies. IT reset password and I modified the lockout policies.
65
2017062813:00Cleaned fluoro and trans
66
201706278:00Intake pipes cleaned resulting in noticable change in FLUORO data. Recommended to maintenance staff intakes cleaning every 6 months.
67
2017062214:45Cleaned fluoro and trans
68
2017060713:00Cleaned fluoro, trans, pH, CTD, pCO2, and optode. Cleared biofouling from tubing between optode and CTD. Flow still seems to be restricted somewhere.
69
2017053014:00cleaned fluoro and trans.
70
201705171000Cleaned fluoro and trans. Deployed/calibrated new pH probe.
71
201705119:30Cleaned fluoro and trans
72
201705031000Cleaned pCO2, CTD, fluoro, trans, optode
73
201704261100Cleaned fluoro and trans
74
20170420Cleaned fluoro and trans
75
20170413Backup CTD SBE19 (2216) deployed
76
201704138:00cleaned trans and fluoro
77
2017032914:30Cleaned CTD, optode, fluoro, trans, pCO2. Added MBARI, experimental sensors for side-by-side comparison. MBARI scientist took verification samples for pH, DIC, DO.
78
2017032215:00Cleaned fluoro and trans
79
20170322-20170323Network device (Lantronix UDS2100) migrated over to a new VLAN. Down for configuration.
80
2017031514:00Cleaned fluoro and trans.
81
2017030813:00Cleaned fluoro and transmissometer.. Redeployed pCO2 sensor, but still working on remote server.
82
2017030110:30Cleaned fluoro and transmissometer..
83
2017022311:00Cleaned fluoro and transmissometer. Installed flowmeter. Will keep flow at 2 L/Min per C-Sense manufacturer request.
84
201702154:30Cleaned trans
85
201702071100Cleaned fluoro and trans
86
20170121 - ongoingData spikes caused by hourly backflushing of the sand filters. Major swell event on 2017-01-25 could have damaged or covered intake opening, causing pumps to draw in sediment.I dove on the intakes and found no damage. Highly turbid water coming from the river is fouling up the sand filters. Problem resolved itself over the course of several days.
87
201701259:00Cleaned fluor and trans. Data seem spikey. Keeping an eye on it.
88
20170120-23Data aquisition services merged with pubdata server. Implemented realtime QARTOD checking.
89
2017011110:00Cleaned fluoro and trans
90
201701048:30Cleaned CTD, optode, fluoro, trans. Removed fouling from line. Removed pH sensor from service.
91
2016120910:30Cleaned fluoro and trans.
92
201612068:45pH probe began reporting squirrelly data, then flatlined.
93
201611288:00Cleaned fluoro and trans
94
201611228:30Cleaned fluoro and trans
95
2016111416:00Cleaned optode, fluoro, trans, pH. Heavy biofouling (mussells, barnacles) throughout system lines was slowing flow. Cleaned out the improved flow.
96
201610315:00pCO2 sensor began reporting bad data
97
201610289:00cleaned fluoro, trans
98
2016092712:00Cleaned fluoro, trans, pH, CTD, and optode
99
20160925-27Server memory full, no data collected.
100
201609157:40
Loading...
Main menu